OpenStudy (anonymous):

y varies inversely as twice x. When x = 3, y = 8. Find y when x = 2.

OpenStudy (lgbasallote):

y varies inversely as twice x so \[y = \frac{k}{2x}\] when x = 3 then y = 8 \[8 = \frac{k}{2(3)} \implies 8 = \frac{k}{6} \implies 48 = k\] find y when x= 2 \[y = \frac{k}{2x}\] \[y = \frac{48}{2(2)}\] \[y = \frac{48}{4}\] \[y = 12\] do you get that?